OpenStudy (anonymous):

cos(x/3)cos(x/3)=1/2[1+cos(2x/3)] is this true or false?

OpenStudy (mathmale):

Have you a table of trig identities? Hint: replace every instance of x/3 with y. Do you then see a familiar trig identity?

OpenStudy (anonymous):

yes its part of the product to sum identities right?

OpenStudy (mathmale):

Look again, please. Hint: cos y cos y = ?

OpenStudy (anonymous):

I still don't understand

OpenStudy (mathmale):

Look in your table of trig identities for an identity for (cos x)^2. There's also an identity for (sin x)^2 that has the same form except for sign.

OpenStudy (anonymous):

ohhh cos^2x=1+cos2x/2 right?

OpenStudy (mathmale):

cos^2x=1+cos2x/2 would be fine with parentheses, not so fine without: cos^2x=(1+cos2x)/2 right?

OpenStudy (anonymous):

yes, so than the answer would be false. correct?

OpenStudy (mathmale):

If you're going to do a lot through OpenStudy, try learning Equation Editor. It's a gem! Explain how you arrived at that conclusion, please.

OpenStudy (mathmale):

\[\cos^2x=\frac{ 1 }{ 2 }(1+\cos 2x)\]

OpenStudy (mathmale):

is an example of how much clearer this identity is when done in Equation Editor. Why do you feel that the original statement is false?

OpenStudy (anonymous):

If the original is \[\cos2x=1/2(1+\cos2x)\] than dividing by 3 would make it false right?

OpenStudy (anonymous):

sorry meant to put the cos squared not 2x

OpenStudy (mathmale):

I see. All the more reason to learn to use Equation Editor. Bet you'll like it once you've tried it. OK. Now I'd agree with your response.

OpenStudy (anonymous):

So i was correct with the statement that the answer is false?